How many grams of dry NH4Cl need to be added to 2.40 L of a 0.400 M solution of ammonia, NH3, to prepare a buffer solution that has a pH of 8.78? Kb for ammonia is 1.8 x 10^-5.
Answer #1

pOH 14.-8.78 5.22 pKb = 4.74 5.22-4.74 = 0.48 NH4 0.400 10048- 3.01 [NH,*) = 0.400 x 3.01=1.207 M moles [NH4-] = 1.207×2.40 L